If %errors of measurement in velocity and mass2%and3%respectively what is the %error in kinetic energy

Grade:11

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
we know KE = 1/2 mv^2
delta(KE)=delta(m)+2delta(v)
now given that,delta(m)=2% =2/100=0.02
and delta(v)=3% =3/100=0.03
so delta(KE)=0.02+2*(0.03)=0.08=8%
so percentage error in Kinetic Energy is 8%. [Ans]
Regards,
Thank you,
Soumya Ranjan Mohanty.